# Scalar-Induced Gravitational Waves from self-resonant preheating in $\bm{\alpha}$-attractor models

### v1

@@ 2 Scalar perturbations
where a tilde means derivation with respect to conformal time $\eta$, related to the cosmic time as $\mathrm{d}\eta=\mathrm{d}t/a$, $\epsilon_{H}=1/2(\phi^{\prime}/\mathcal{H})^{2}$ is the first slow-roll parameter and $\mathcal{H}=a^{\prime}/a$ is the conformal Hubble rate. Here, $v_{\bm{k}}=a\left[\delta\phi_{\bm{k}}+\phi^{\prime}\Phi_{\bm{k}}/\mathcal{H}\right]$ is the so-called MS variable, a combination of the field and metric perturbations, $\delta\phi_{\bm{k}}$ and $\Phi_{\bm{k}}$, respectively. See Mukhanov:1990me ; Baumann:2009ds for details. As shown in del-Corral:2024vcm , during a self-resonant phase, one can perturbatively solve the evolution equation of the inflation field as a series expansion of a small parameter $\beta$ as follows
@@
whose time-evolution is shown in Fig. 4 for the analytical model of del-Corral:2024vcm (dashed green line), along with the full numerical solution (solid black line). Notably, as $\alpha$ decreases, the curvature perturbation $\mathcal{R}_{\bm{k}}$ amplifies during the first e-folds after the end of inflation, caused by the self-resonance effect.
This amplification is crucial for scalar-induced gravitational waves in the second order. Note that we study all the modes within the instability band (See Fig. 1), which constitute those that exit the horizon during inflation as well as those that never exit the horizon but fall into the instability band with wavelengths greater than the Jeans length $R_{J}\simeq\sqrt{2\langle c_{s}\rangle^{2}/3H^{2}}$ del-Corral:2023apl ; draft-alpha-PBH where
@@
[…]
@@
where $\lambda=-\frac{2M^{2}}{9\alpha}<0$ for the T-model and $\lambda=\frac{7M^{2}}{9\alpha}>0$ for the E-model. It was shown in del-Corral:2023apl ; draft-alpha-PBH that both the modes that exit and those that never exit the horizon during inflation could end up inside the instability band with identical behavior. However, the modes that never exit the horizon during inflation and enter the instability band get amplified largely and quickly. As we can see in Fig. 4, the amplification is significant (a few orders of magnitude higher) for the modes with $k>k_{\rm end}$ compared to those with $k<k_{\rm end}$, where $k_{\rm end}$ is the wave number of the mode that exits the horizon at the end of inflation. This amplification of curvature perturbations is essential for computing the SIGWs, which we present in the next section.
Figure 4: Evolution of the curvature perturbation $\mathcal{R}_{\bm{k}}$ after the end of inflation for (a) a T-model with $\alpha=10^{-3}$ and (b) an E-model with $\alpha=10^{-3}$. The evaluation is carried out for 5 e-folds after the end of inflation. Numerical computation is shown as a solid black line, and the analytical model as a dashed green line. The vertical grey line marks the scale that exits the horizon at the end of inflation, $k_{\text{end}}$. For comparison, black dotted curves show the numerical computation of curvature perturbations at the end of inflation. Figure adapted from del-Corral:2024vcm .
3 Tensor perturbations
After computing the scalar perturbations during preheating, we aim to analyze the tensor perturbations to determine whether the amplification of scalar perturbations leaves an imprint on the power spectrum of SIGWs. We first use perturbation theory in Sec. 3.1 to estimate the energy density of gravitational waves and then, in Sec. 3.2, we explain the numerical strategy followed to achieve that goal.
3.1 Energy density of gravitational waves
Going to the 3rd order perturbation of the action (4) in scalar ($\Phi,\,\delta\phi$) and tensor ($h_{ij}^{\lambda}$) fluctuations, we obtain the second-order [^1] At the level of first order perturbed Einstein equation (or at the level of second order action) scalar and tensor modes do not couple, whereas the perturbed Einstein equations at the second-order (i.e., at the level of 3rd order perturbed action) we can have scalar modes acting as additional source to the generation of gravitational waves on top of the background as (12) illustrates. equations of motion for the tensor modes $h_{\textbf{k}}$ in Fourier space as Domenech:2021 ; Baumann:2007 ; Ananda:2006 ; Assadullahi:2009
@@ 3.1 Energy density of gravitational waves
$\mathcal{S}(\bm{k},t)=\int\frac{\mathrm{d}^{3}\tilde{k}}{(2\pi)^{3/2}}\tilde{k}^{2}(1-\mu^{2})\left(4\Phi_{\tilde{\bm{k}}}\Phi_{\bm{k}-\tilde{\bm{k}}}+2\delta\phi_{\tilde{\bm{k}}}\delta\phi_{\bm{k}-\tilde{\bm{k}}}\right).$ (13)
The above source term contribution originates from the scalar-scalar-tensor vertex interaction term in the third-order perturbed expansion of the action (4). Here, we neglect 4th and higher-order contributions on the expectation that any further scalar self-interactions are negligible. As we can see from Fig. 4, the curvature perturbation, which is a combination of metric and scalar field fluctuation, remains smaller than unity even in the regime of self-resonance amplification during the preheating phase. If we consider the scales of full non-linear regime, then the effects of non-Gaussianities are important Papanikolaou:2024kjb ; He:2024luf , which we defer for future investigations. Since the source term (13) on the right-hand side of (12) reflects the fact that these GWs are no longer free-propagating waves but rather a metric distortion arising from terms quadratic in the scalar perturbations Assadullahi:2009 , they are called SIGWs. Moreover, the integral over $\tilde{k}$ reflects that the computation of SIGWs involves a convolution of different modes, meaning the contribution from any individual mode is diluted and mixed with contributions from other modes. Consequently, the amplification of the gravitational waves arises from the collective contribution of all modes within the Hubble horizon. In our case, the significant contribution to the amplification comes from the modes $k>k_{\rm end}$ (See Fig. 4). In App. A we show the derivation of the source term $\mathcal{S}(\bm{k},t)$ from the perturbed Einstein equations and how to express it in terms of the metric scalar perturbations, as shown in eqn. (13).
We can compute the energy density of GWs as the 00 component of the pseudo-energy momentum tensor of GWs Domenech:2021 , that is
$\rho_{\text{GW}}=M_{\mathrm{Pl}}^{2}\int\mathrm{d}\ln{k}\frac{k^{3}}{16\pi^{2}}\Bigg{\{}\sum_{\lambda}\langle\dot{h}_{\bm{k}}^{(\lambda)}\dot{h}_{-\bm{k}}^{(\lambda)}\rangle^{!}+\frac{k^{2}}{a^{2}}\langle h_{\bm{k}}^{(\lambda)}h_{-\bm{k}}^{(\lambda)}\rangle^{!}\Bigg{\}},$ (14)
@@
$\begin{split}\mathcal{P}_{h}^{\text{SI}}(k,t)=\frac{16g^{2}(k,t)}{k}\int_{0}^{\infty}\mathrm{d}\tilde{k}\int_{-1}^{1}\mathrm{d}\mu\,\frac{\tilde{k}^{3}\,(1-\mu^{2})^{2}}{|k-\tilde{k}|^{3}}\mathcal{P}_{\Phi}(\tilde{k})\mathcal{P}_{\Phi}(|k-\tilde{k}|),\end{split}$ (16)
where SI stands for scalar-induced, $\mu$ is the cosine of the angle between ingoing and outgoing scales, and $\mathcal{P}_{\Phi}(k)$ is the power spectrum of the scalar perturbations $\Phi_{\bm{k}}$. It is defined similarly to (15), but a Heaviside function is introduced in this case to avoid the contribution to the integral coming from the non-linear regime of $\mathcal{P}_{\Phi}$ Assadullahi:2009 , that is
@@
where the function $g(k,t)$ is defined in eqn. (46). App. B shows the details about the derivation of (16) following the standard derivations of GWs production during matter-dominated scenarios as well as some analytical estimates for the spectral energy density of GWs. The latter, labeled as $\Omega_{GW}$, is often used instead of $\rho_{GW}$ to compare theoretical predictions with current constraints and future observations. It represents the energy density per logarithm of $k$ over the critical density and is given by
@@
where we have used $a(t_{\text{rh}})H(t_{\text{rh}})=k_{L}$. After the end of preheating, the fractional energy density redshifts at sub-Hubble scales as any non-interacting relativistic particles. Therefore, the present energy density of gravitational waves is approximated by Assadullahi:2009
$\begin{split}\Omega_{\text{GW}}^{\text{SI}}(k,t_{0})&\simeq\Omega_{\gamma}^{0}\,\Omega_{\text{GW}}^{\text{SI}}(k,t_{\text{rh}})\\ &\simeq\frac{4\,\Omega_{\gamma}^{0}}{3}\left(\frac{k}{k_{L}}\right)^{2}\frac{g^{2}(k,t_{\text{rh}})}{k}\int_{0}^{\infty}\mathrm{d}\tilde{k}\int_{-1}^{1}\mathrm{d}\mu\,\frac{\tilde{k}^{3}\,(1-\mu^{2})^{2}}{|k-\tilde{k}|^{3}}\,\mathcal{P}_{\Phi}(\tilde{k})\mathcal{P}_{\Phi}(|k-\tilde{k}|),\end{split}$ (21)
where $\Omega_{\gamma}^{0}\simeq 1.2\times 10^{-5}$ is the present energy density of photons, $t_{0}$ represents the present epoch, and we have substituted in the last line the power spectrum from eqn. (16), see Assadullahi:2009 for details. This quantity is usually expressed as a function of the GW frequency $f$ rather than the wave number $k$. The relationship between them is given by Frosina:2023nxu
@@
[…]
where $a_{k}$ is the scale factor at the time the mode $k$ enters the horizon, defined as $t_{k}$. The factor $1/2$ accounts for averaging over harmonic oscillations of modes deep inside the horizon Figueroa:2019paj . The background tensor power spectrum evolves as follows after inflation:
@@
If the mode enters during preheating, a suppression effect exists due to the $a^{4}H^{2}\propto a$ term in eqn. (28). This means that the fractional energy density gets suppressed as $\Omega^{\text{B}}_{\text{GW}}\propto a_{k}/a$ for the modes entering during preheating Figueroa:2019paj ; Caprini:2018mtu ; Sahni:2001qp ; Sahni:1990tx ; Mishra:2021wkm , that is, for $k>k_{L}$. This effect disappears when reheating begins. At this point, and as in the scalar-induced case, the fractional energy density redshifts as any non-relativistic particles. This implies we can use the following parametrization for the fractional energy density of BGWs evaluated at the present time:
@@
3.2 Numerical strategy
To numerically compute the total fractional energy density of SIGWs, we begin by computing the time evolution of the curvature perturbations $\mathcal{R}_{\bm{k}}$, following the method outlined in Sec. 2. The results are stored in $k$-space. Next, the power spectrum of the curvature perturbations, $\mathcal{P}_{\mathcal{R}}$, is computed and evaluated at the end of the short self-resonance stage, as it remains constant for the modes of interest thereafter. At this stage, the universe behaves as matter-dominated, allowing us to use the following relation between $\mathcal{P}_{\mathcal{R}}$ and $\mathcal{P}_{\Phi}$ Assadullahi:2009 ; Baumann:2007
$\mathcal{P}_{\mathcal{R}}(k)=\frac{9}{25}\mathcal{P}_{\Phi}(k),$ (31)
where $\mathcal{P}_{\mathcal{R}}$ is defined analogously to (17). This relation is then substituted into (21). Since the curvature perturbations can only be obtained numerically in discrete steps of $k$, the integral over $\tilde{k}$ in (21) is discretized. Specifically, the $k$-space integral is divided into 100 uniform points per logarithmic interval of $k$. This procedure is repeated for each value of $k$ in $k\in[k_{L},k_{U}]$ and evaluated at the end of the preheating stage, $t_{\text{rh}}$, which we consider to last approximately 5 e-folds.
@@ 3.2 Numerical strategy
Fig. 5 shows the computation of the total fractional energy density of GWs for the two models under consideration (T-model as dashed lines and E-model as solid lines) and for different values of the parameter $\alpha$. We have also included the Starobinsky model Starobinsky:1980te , which corresponds to an E-model with $\alpha=1$, for comparison purposes. An interesting feature of Fig. 5 is that the amplification peak for high-$k$ in the scalar perturbations of Fig. 4 is also present in the energy density of GWs for small $\alpha$, since $\Omega_{\text{GW}}^{\text{TOT}}\sim\Omega_{\text{GW}}^{\text{SI}}$. However, for higher values of $\alpha$, such as the Starobinsky case, the amplification of scalar perturbations is negligible and thus $\Omega_{\text{GW}}^{\text{TOT}}\sim\Omega_{\text{GW}}^{\text{B}}$. Nonetheless, due to the convolution mentioned earlier, this contribution from the peak gets diluted among the rest of the modes inside the horizon, and thus the peak in $\Omega_{\text{GW}}^{\text{TOT}}$ is not as sharp as it is in $\mathcal{P}_{\mathcal{R}}$, which makes it difficult to track back the initial feature in the power spectrum causing the amplification. In App. B we show some analytical approximations that explain the shape of $\Omega_{\text{GW}}^{\text{SI}}$. We also remark here that, as anticipated, decreasing $\alpha$ leads to a growth of the metric perturbations, which enhances gravitational wave production in the very high-frequency (VHF) band. Although there is no current GW detector for the VHF band, we can use the BBN bound to restrict the model’s parameters. This proceeds as follows. The GWs at the onset of the BBN contribute to the radiation density and can potentially change the expansion rate of the universe, and thus the abundance of the light elements. Measurements of this abundance lead to an equivalent number of additional relativistic species of 1.4 neutrino degrees of freedom, which translates into the following upper bound for the total fractional energy density of GWs Smith:2006nka ; Maggiore:1999vm
@@
Figure 6: Integrated fractional energy density, eqn. (21), normalized to $\mathcal{I}_{\text{BBN}}$ as a function of $\alpha$ for the T- and E-models.
@@
This bound is represented as the horizontal red band in Fig. 5, which is reached for the E-model with $\alpha=10^{-3.25}$. It is important to remark that the bound is on the integral over the whole spectrum of modes, which implies that $\Omega_{\text{GW}}^{\text{TOT}}$ could present a peak above $\mathcal{I}_{\text{BBN}}$, provided it is narrow so that $\mathcal{I}_{\Omega}<\mathcal{I}_{\text{BBN}}$. Thus, Fig. 5, is not good enough to establish a lower bound on $\alpha$. Instead, Fig. 6 shows the result of the numerical integration of the fractional energy density (21) as a function of $\alpha$ for the T- and E-models, normalized to $\mathcal{I}_{\text{BBN}}$. Here, we clearly see a limiting value of $\alpha$, for which the BBN bound is reached (red shaded area). This corresponds to $\log_{10}(\alpha)\simeq-3.54$ for the T-model and to $\log_{10}(\alpha)\simeq-3.17$ for the E-model. Since the amplification occurs rapidly ($\mathcal{O}(1)$ e-fold), this essentially restricts the duration of preheating. Also, we can use (1) to put a lower bound on the tensor-to-scalar ratio, that is
$r_{T}>9.61\times 10^{-7}\qquad\text{and}\qquad r_{E}>2.25\times 10^{-6},$ (33)
where the suffixes T and E stand for the different bounds on $\alpha$ for each model. The lower boundon $r_{E}$ rests on the fact that the E-model potential is asymmetric and thus self-resonance is stronger in this case, for the same value of $\alpha$, due to the drift terms in the potential expansion del-Corral:2024vcm . This lower bound, although far from the sensitivity of the future satellite observer LiteBIRD ($r\sim 1.0\times 10^{-3}$) LiteBIRD:2022nob ; LiteBIRD:2023aov , agrees with the bound using BICEP2 and Planck data Tristram:2021tvh .

### v2

@@ II Scalar perturbations
where a tilde means derivation with respect to conformal time $\eta$, related to the cosmic time as $\mathrm{d}\eta=\mathrm{d}t/a$. Also,
$\epsilon_{H}=-\frac{\dot{H}}{H^{2}}=\frac{1}{2}\left(\frac{\phi^{\prime}}{\mathcal{H}}\right)^{2},\qquad\mathcal{H}=\frac{a^{\prime}}{a},$ (6)
are the first slow-roll parameter and the conformal Hubble rate, respectively, and
$v_{\bm{k}}=a\left[\delta\phi_{\bm{k}}+\frac{\phi^{\prime}\Phi_{\bm{k}}}{\mathcal{H}}\right]$ (7)
is the so-called MS variable, a combination of the field and metric perturbations, $\delta\phi_{\bm{k}}$ and $\Phi_{\bm{k}}$, respectively. See [66, 67] for details. As shown in [23], during a self-resonant phase, one can perturbatively solve the evolution equation of the inflation field as a series expansion of a small parameter $\beta$ as follows
@@
whose time-evolution is shown in Fig. 4. We observe that, as $\alpha$ decreases, the curvature perturbation $\mathcal{R}_{\bm{k}}$ rapidly amplifies during the first e-folds after the end of inflation, caused by the self-resonance effect. This is essentially what characterizes the robustness of our result. Even a short duration of preheating of 1-2 e-folds will produce this amplification for the perturbations. Fig. 4 shows the curvature perturbations for a T- and an E-model with the values of $\alpha$ corresponding to the lower bounds derived in this work. The evaluation is performed from the end of inflation to 2 e-folds after, in steps of 0.1 e-folds. One can observe that the amplification is significant even for a relatively short duration of preheating, which confirms our statement. If one extends the evaluation of this plot further than 2 e-folds after the end of inflation, the curvature perturbations would not increase, as the self-resonance effects are only present during the first e-folds after inflation.
Figure 4: Curvature perturbation $\mathcal{R}_{\bm{k}}$ as a function of the comoving wavenumber $k$ for (a) T-model and (b) E-model. The evaluation is performed from the end of inflation to 2 e-folds after in steps of 0.1 e-folds. The vertical dashed lines mark the scale $k_{\text{end}}$.
Using the perturbed Einstein equations in the Newtonian gauge, one can obtain the following set of equations that allows us to obtain the metric and the field perturbations
$\mathcal{R}_{\bm{k}}=\frac{H^{-1}\dot{\Phi}_{\bm{k}}+\Phi_{\bm{k}}}{\epsilon_{H}}+\Phi_{\bm{k}},$ (11a)
$\delta\phi_{\bm{k}}=\sqrt{2\epsilon_{H}}\left(\mathcal{R}_{\bm{k}}-\Phi_{\bm{k}}\right).$ (11b)
These perturbations also carry the self-resonance effects from the curvature perturbation and, as we show below, are essential for the computation of the SIGWs.
This amplification is crucial for scalar-induced gravitational waves in the second order. Note that we study all the modes within the instability band (See Fig. 1), which constitute those that exit the horizon during inflation as well as those that never exit the horizon but fall into the instability band with wavelengths greater than the Jeans length $R_{J}\simeq\sqrt{2\langle c_{s}\rangle^{2}/3H^{2}}$ [22, 44] where
@@
where $\langle\dots\rangle$ means averaging over a time period larger than the oscillation period but smaller than the Hubble time [68]. The potential $V(\phi)$ in (12) is the inflaton potential. For the E-model and T-model alpha-attractor potentials (2), the analytical estimate for the average sound speed is
@@
where $\lambda=-\frac{2M^{2}}{9\alpha}<0$ for the T-model and $\lambda=\frac{7M^{2}}{9\alpha}>0$ for the E-model. It was shown in [22, 44] that both the modes that exit and those that never exit the horizon during inflation could end up inside the instability band with identical behavior. However, the modes that never exit the horizon during inflation and enter the instability band get amplified largely and quickly. As we can see in Fig. 4, the amplification is significant (a few orders of magnitude higher) for the modes with $k>k_{\rm end}$ compared to those with $k<k_{\rm end}$, where $k_{\rm end}$ is the wave number of the mode that exits the horizon at the end of inflation, marked by the vertical dashed lines. This amplification of curvature perturbations is essential for computing the SIGWs, which we present in the next section.
III Tensor perturbations
After computing the scalar perturbations during preheating, we aim to analyze the tensor perturbations to determine whether the amplification of scalar perturbations leaves an imprint on the power spectrum of SIGWs. We first use perturbation theory in Sec. III.1 to estimate the energy density of gravitational waves and then, in Sec. III.2, we explain the numerical strategy followed to achieve that goal.
III.1 Energy density of gravitational waves
Going to the 3rd order perturbation of the action (3) in scalar ($\Phi,\,\delta\phi$) and tensor ($h_{ij}^{\lambda}$) fluctuations, we obtain the second-order  [^1] At the level of first order perturbed Einstein equation (or at the level of second order action) scalar and tensor modes do not couple, whereas the perturbed Einstein equations at the second-order (i.e., at the level of 3rd order perturbed action) we can have scalar modes acting as additional source to the generation of gravitational waves on top of the background as (14) illustrates. equations of motion for the tensor modes $h_{\textbf{k}}$ in Fourier space as [41, 38, 36, 37]
@@ III.1 Energy density of gravitational waves
[…]
The above source term contribution originates from the scalar-scalar-tensor vertex interaction term in the third-order perturbed expansion of the action (3), which is
$\delta^{(3)}S=\int d^{3}x\,dt\Bigg\{-a\,h^{ij}\,\partial_{i}\Phi\,\partial_{j}\Phi+\frac{a}{2}\,h^{ij}\,\partial_{i}\delta\varphi\,\partial_{j}\delta\varphi\Bigg\}.$ (16)
Here, we neglect 4th and higher-order contributions on the expectation that any further scalar self-interactions are negligible. It is important to emphasize that the source term in Eq. (15) is expressed in terms of the gauge-invariant variables $\Phi_{\bm{k}}$ and $\delta\phi_{\bm{k}}$ [66], as a result, no gauge ambiguities are expected to arise. As we can see from Fig. 4, the curvature perturbation, which is a combination of metric and scalar field fluctuation, remains smaller than unity even in the regime of self-resonance amplification during the preheating phase. If we consider the scales of the full non-linear regime, then the effects of non-Gaussianities are important [69, 70], which we defer for future investigations. Since the source term (15) on the right-hand side of (14) reflects the fact that these GWs are no longer free-propagating waves but rather a metric distortion arising from terms quadratic in the scalar perturbations [37], they are called SIGWs. Moreover, the integral over $\tilde{k}$ reflects that the computation of SIGWs involves a convolution of different modes, meaning the contribution from any individual mode is diluted and mixed with contributions from other modes. Consequently, the amplification of the gravitational waves arises from the collective contribution of all modes within the Hubble horizon. In our case, the significant contribution to the amplification comes from the modes $k>k_{\rm end}$ (See Fig. 4). In App. A we show the derivation of the source term $\mathcal{S}(\bm{k},t)$ from the perturbed Einstein equations and how to express it in terms of the metric scalar perturbations, as shown in eqn. (15).
We can compute the energy density of GWs as the 00 component of the pseudo-energy momentum tensor of GWs [41], that is
$\begin{split}\rho_{\text{GW}}&=M_{\mathrm{Pl}}^{2}\int\mathrm{d}\ln{k}\frac{k^{3}}{16\pi^{2}}\\ &\times\Bigg\{\sum_{\lambda}\langle\dot{h}_{\bm{k}}^{(\lambda)}\dot{h}_{-\bm{k}}^{(\lambda)}\rangle^{!}+\frac{k^{2}}{a^{2}}\langle h_{\bm{k}}^{(\lambda)}h_{-\bm{k}}^{(\lambda)}\rangle^{!}\Bigg\},\end{split}$ (17)
@@
[…]
where SI stands for scalar-induced, $\mu$ is the cosine of the angle between ingoing and outgoing scales, and $\mathcal{P}_{\Phi}(k)$ and $\mathcal{P}_{\Phi}(k)$ are the power spectra of the scalar $\Phi_{\bm{k}}$ and field $\delta\phi_{\bm{k}}$ perturbations. They are defined similarly to (18), but a Heaviside function is introduced in this case to avoid the contribution to the integral coming from the non-linear regime [37], that is
@@
where the function $g(k,t)$ is defined in eqn. (51), and the same occurs for the field perturbations. App. B shows the details about the derivation of (19) following the standard derivations of GWs production during matter-dominated scenarios as well as some analytical estimates for the spectral energy density of GWs. The latter, labeled as $\Omega_{\text{GW}}$, is often used instead of $\rho_{\text{GW}}$ to compare theoretical predictions with current constraints and future observations. It represents the energy density per logarithm of $k$ over the critical density and is given by
@@
where we have used $a(t_{\text{rh}})H(t_{\text{rh}})=k_{L}$. After the end of preheating, the fractional energy density redshifts at sub-Hubble scales as any non-interacting relativistic particles. Therefore, the present energy density of gravitational waves is approximated by [37]
$\begin{split}\Omega_{\text{GW}}^{\text{SI}}(k,t_{0})&\simeq\frac{4\,\Omega_{\gamma}^{0}}{3}\left(\frac{k}{k_{L}}\right)^{2}\frac{g^{2}(k,t_{\text{rh}})}{k}\int_{0}^{\infty}\mathrm{d}\tilde{k}\int_{-1}^{1}\mathrm{d}\mu\,\frac{\tilde{k}^{3}(1-\mu^{2})^{2}}{|k-\tilde{k}|^{3}}\bigg[\mathcal{P}_{\Phi}(\tilde{k})\mathcal{P}_{\Phi}(|k-\tilde{k}|)+\frac{1}{2}\mathcal{P}_{\delta\phi_{\bm{k}}}(\tilde{k})\mathcal{P}_{\delta\phi_{\bm{k}}}(|k-\tilde{k}|)\bigg]\\ &\simeq\Omega_{\text{GW},\Phi}^{\text{SI}}(k,t_{\text{rh}})+\Omega_{\text{GW},\delta\phi}^{\text{SI}}(k,t_{\text{rh}}).\end{split}$ (24)
where $\Omega_{\gamma}^{0}\simeq 1.2\times 10^{-5}$ is the present energy density of photons, $t_{0}$ represents the present epoch, and we have substituted in the last line the power spectrum from eqn. (19), see [37] for details. This quantity is usually expressed as a function of the GW frequency $f$ rather than the wave number $k$. The relationship between them is given by [71]
@@
$\mathcal{T}^{\,2}_{h}(k,t)=\begin{cases}1&t<t_{k}\\ \frac{1}{2}\left(\frac{a_{k}}{a}\right)^{2}&t>t_{k}\end{cases}$ (29)
where $a_{k}$ is the scale factor at the time the mode $k$ enters the horizon, defined as $t_{k}$. The factor $1/2$ accounts for averaging over harmonic oscillations of modes deep inside the horizon [72]. The background tensor power spectrum evolves as follows after inflation:
@@
If the mode enters during preheating, a suppression effect exists due to the $a^{4}H^{2}\propto a$ term in eqn. (31). This means that the fractional energy density gets suppressed as $\Omega^{\text{B}}_{\text{GW}}\propto a_{k}/a$ for the modes entering during preheating [72, 73, 74, 75, 76], that is, for $k>k_{L}$. This effect disappears when reheating begins. At this point, and as in the scalar-induced case, the fractional energy density redshifts as any non-relativistic particles. This implies we can use the following parametrization for the fractional energy density of BGWs evaluated at the present time:
@@
III.2 Numerical strategy
To numerically compute the total fractional energy density of SIGWs, we begin by computing the time evolution of the curvature perturbations $\mathcal{R}_{\bm{k}}$, following the method outlined in Sec. II. Then, using eqns. (11a) and (11b) we compute the metric and field perturbations $\Phi_{\bm{k}}$ and $\delta\phi_{\bm{k}}$. The results are stored in $k$-space. Next, the power spectrum of the metric and field perturbations, $\mathcal{P}_{\Phi}$ and $\mathcal{P}_{\delta\phi}$, are computed and evaluated at the end of the short self-resonance stage, as it remains constant for the modes of interest thereafter. Since the scalar perturbations can only be obtained numerically in discrete steps of $k$, the integral over $\tilde{k}$ in (24) is discretized. Specifically, the $k$-space integral is divided into 100 uniform points per logarithmic interval of $k$. This procedure is repeated for each value of $k$ in $k\in[k_{L},k_{U}]$ and evaluated at the end of the preheating stage, $t_{\text{rh}}$, which we consider to range from 1 to 5 e-folds after the end of inflation.
@@ III.2 Numerical strategy
Fig. 5 shows the computation of the total fractional energy density of GWs for the two models under consideration (T-model as dashed lines and E-model as solid lines) and for different values of the parameter $\alpha$. The length of preheating is chosen as 5 e-folds in this case. We have also included the Starobinsky model [77], which corresponds to an E-model with $\alpha=1$, for comparison purposes. An interesting feature of Fig. 5 is that the amplification peak for high-$k$ in the scalar perturbations of Fig. 4 is also present in the energy density of GWs for small $\alpha$, since $\Omega_{\text{GW}}^{\text{TOT}}\sim\Omega_{\text{GW}}^{\text{SI}}$. However, for higher values of $\alpha$, such as the Starobinsky case, the amplification of scalar perturbations is negligible and thus $\Omega_{\text{GW}}^{\text{TOT}}\sim\Omega_{\text{GW}}^{\text{B}}$. Nonetheless, due to the convolution mentioned earlier, this contribution from the peak gets diluted among the rest of the modes inside the horizon, and thus the peak in $\Omega_{\text{GW}}^{\text{TOT}}$ is not as sharp as it is in $\mathcal{P}_{\Phi}$ and $\mathcal{P}_{\delta\phi}$, which makes it difficult to track back the initial feature in the power spectrum causing the amplification. In App. B we show some analytical approximations that explain the shape of $\Omega_{\text{GW}}^{\text{SI}}$. We also remark here that, as anticipated, decreasing $\alpha$ leads to a growth of the metric perturbations, which enhances gravitational wave production in the very high-frequency (VHF) band. Although there is no current GW detector for the VHF band, we can use the BBN bound to restrict the model’s parameters. This proceeds as follows. The GWs at the onset of the BBN contribute to the radiation density and can potentially change the expansion rate of the universe, and thus the abundance of the light elements. Measurements of this abundance lead to an equivalent number of additional relativistic species of 1.4 neutrino degrees of freedom, which translates into the following upper bound for the total fractional energy density of GWs [52, 53]
@@
Figure 6: Integrated fractional energy density, eqn. (24), normalized to $\mathcal{I}_{\text{BBN}}$ as a function of $\alpha$ and the duration of preheating ($\Delta N$) for the T- and E-models. One finds that extending preheating beyond 2 e-folds does not affect the moment when the BBN bound is reached. The reason is that after two e-folds all relevant (amplified) perturbations have entered the horizon, making their contribution to the integral (34) negligible.
@@
This bound is represented as the horizontal red band in Fig. 5, which is reached for the E-model with $\alpha=10^{-3.25}$. It is important to remark that the bound is on the integral over the whole spectrum of modes, which implies that $\Omega_{\text{GW}}^{\text{TOT}}$ could present a peak above $\mathcal{I}_{\text{BBN}}$, provided it is narrow so that $\mathcal{I}_{\Omega}<\mathcal{I}_{\text{BBN}}$. Thus, Fig. 5, is not good enough to establish a lower bound on $\alpha$. Instead, Fig. 6 shows the result of the numerical integration of the fractional energy density (24) as a function of $\alpha$ and the duration of preheating ($\Delta N$) for the T- and E-models, normalized to $\mathcal{I}_{\text{BBN}}$. Here, we clearly see a limiting value of $\alpha$, that we call $\alpha_{\text{BBN}}$, for which the BBN bound (red shaded area) is reached for each $\Delta N$. The results can be seen in Table 1. Since the amplification occurs rapidly ($\mathcal{O}(1)$ e-fold), this essentially restricts the duration of preheating. Also, we can use (1) to put a lower bound on the tensor-to-scalar ratio , which can be seen in Table 1 as $r_{\text{BBN}}$. In general, one can observe that $\alpha_{\text{BBN}}$ is higher for the E-model compared to the T-model for the same $\Delta N$. This rests on the fact that the E-model potential is asymmetric and thus self-resonance is stronger in this case, for the same value of $\alpha$, due to the drift terms in the potential expansion [23]. This lower bounds, although far from the sensitivity of the future satellite observer LiteBIRD ($r\sim 1.0\times 10^{-3}$) [78, 79], agrees with the bound using BICEP2 and Planck data [14]. This bound can be compared with the results of [80] where the authors have obtained a bound on the tensor-to-scalar ratio, based on the latest constraints on the running of the scalar spectral index. Although no precise bound for $r$ is given, it is shown that in three-parameter non-trivial slow-roll inflationary models, if $r\lesssim 10^{-4}$, then the running of the spectral index must satisfy $\alpha_{s}<-10^{-3}$, placing it near the threshold of detectability by upcoming experiments. It is crucial to note that the bound in [80] corresponds to non-trivial single-field models that introduce a large running of spectral index due to a specific potential shape (that has a large non-trivial value for the third derivative of the inflaton potential). In contrast, our investigation is about the simplest single field models of universality class [7].
T-Model E-Model
$\Delta N$ $\log(\alpha_{\text{BBN}})$ $r_{\text{BBN}}$ $\log(\alpha_{\text{BBN}})$ $r_{\text{BBN}}$
[…]
Table 1: Lower bound on $\alpha$ and the corresponding tensor-to-scalar ratio $r$ for each duration of preheating $\Delta N$ from 1 to 5 e-folds.